OpenStudy (anonymous):

An initial population of 745 quail increases at an annual rate of 16%. Write an exponential function to model the quail population. What will the approximate population be after 4 years?

OpenStudy (anonymous):

@Mathmuse @rizwan_uet

OpenStudy (anonymous):

is this correct>?? V = a(1 + g)t V = 745(1 + .16)4 V = 745(1.16)4 V ≈ 745(1.81063936) V ≈ 1349

OpenStudy (anonymous):

V = 745(1 + .16)^t

OpenStudy (anonymous):

t=4 i thought?

OpenStudy (anonymous):

@hba @ParthKohli @ma

OpenStudy (anonymous):

yes, you've got it

OpenStudy (anonymous):

obviously t=4 after 4 years

OpenStudy (anonymous):

There are two parts to the question: Write the modelling equation: @stgreen's equation with time left as variable 't' estimate after 4 years: your original answer with 4 subbed in

OpenStudy (anonymous):

^seconded